Solve for x.
log (base 2)[x^2 + 5] = log(base 2)6x

Answer by lwsshak3(6491)

Solve for x.
log (base 2)[x^2 + 5] = log(base 2)6x
..
log2(x^2+5)=log2(6x)
log2(x^2+5)-log2(6x)=0
log2(x^2+5)/(6x)=0
convert to exponential form.
base raised to the logarithm of a number=number. In this case the base=2,the logarithm of the number=0, and the number=(x^2+5)/(6x)
2^0=(x^2+5)/(6x)=1
x^2+5=6x
x^2-6x+5=0
(x-5)(x-1)=0
x=5
x=1